Hello...
I am trying to figure out what q is when 0 = -2q^2 + 820q - 8000
I can see that i need to add 8000 to both sides...
-2q^2 + 820q = 8000
but then i am not sure where to go from here... what are the rules for this? thanks!

Answer by mananth(16946)

0 = -2q^2 + 820q - 8000
divide both sides by -1
0=2q^2-820q+8000
/2
q^2-410+4000=0
...
q^2-400q-10q+4000=0
(split the middle term into two terms such that their sum is -410 and product is 4000)
...
q(q-400)-10(q-400)=0
(q-400)(q-10)=0
q=400 OR q=10